OpenStudy (anonymous):

Factoring Binomials in the Degree of 3 Can someone list the steps when Factoring Binomials in the Degree of 3 or higher please? Example:

OpenStudy (anonymous):

\[r ^{3}s ^{3}+a ^{3}b ^{3}\]

OpenStudy (anonymous):

formula for factoring sum of cubes is a3 + b3 = ( a + b) (a2b - ab2 + b3) so: r3s3 + a3b3 = (rs + ab)(r2s2ab - rsa2b2 + a2b2) consider rs to be a and ab to be b in the formula

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

The factors will be a binomial and a trinomial Since it is the sum of two cubes the first factor will be the sum of its cube root The next factor which is a trinomial will have alternate +, - + terms The first term of which is the square of the cube root of the first term times the cube root of the second term, etc

OpenStudy (anonymous):

Happy to be of help